OpenStudy (anonymous):

m+3n=1 solve for n **Show Work**

OpenStudy (saifoo.khan):

3n = 1 -m n = (1-m)/3

OpenStudy (anonymous):

m+3n=1 subtract m 3n=1-m divide by 3 n=1/3 - m/3